# CLIP-BEVFormer: Enhancing Multi-View Image-Based BEV Detector with Ground Truth Flow

### v1

@@ 1 Introduction
Autonomous Driving Systems (ADS) have witnessed rapid advancements, revolutionizing the landscape of transportation [13, 9, 14, 16, 6]. The current state-of-the-art vision based ADS approaches heavily rely on Bird’s Eye View (BEV) representations extracted from multi-view images to perceive and understand the surrounding environment. The BEV detector, which acts as the backbone of ADS, transforms multi-view images into a top-down view representation known as the BEV feature map [20, 19, 12, 17, 31]. The effectiveness of the BEV detector significantly influences the success of perception tasks in autonomous driving, underscoring its pivotal role in enhancing the system’s situational awareness [17, 14, 16].
@@ 2.2 Vision-Language Models
Vision-language models (VLMs) have shown great promise for learning good representations for variety of downstream tasks [22, 23, 36, 34, 15, 1]. The success of VLMs have been driven by training transformers on large scale image-text pairs data collected from web using contrastive learning  [22, 15]. Among VLMs, CLIP [22] which have been trained on 400 million pairs of data have shown remarkable zero-shot generalization on various image recognition tasks. The contrastive learning mechanism which maps image and text pairs to a joint embedding space has been pivotal for CLIP success. While VLMs generally show good performance on downstream tasks without any fine-tuning, prompting and fine-tuning techniques have also been employed for adaptation of VLMs to a new downstream task  [40, 10].
@@ 4 Experiments
In this section, we present our experiment dataset, metrics and settings in Sec. 4.1. We validate our CLIP-BEVFormer by conducting experiments on 3D perception, long-tail cases, and robustness tasks, as illustrated in Sec. 4.2, Sec. 4.3, and Sec.4.4, respectively. We demonstrate our ablation studies in Sec.4.5. We additionally provide visualization results, comparing CLIP-BEVFormer with baseline models, as presented in Section 4.6.
@@ 4.1 Implementation Details
Dataset. Our experiments are conducted on public nuScenes dataset[3], which is a commonly used dataset in autonomous driving area. It contains 1000 scenes, each with a $\sim$20s duration, and the keyframes are annotated at 2Hz. Each scene is captured with 6 cameras covering the entire 360 ${}^{\circ}$ field-of-view. Overall, the nuScenes dataset contains 1.4M 3D annotated bounding boxes of 10 object categories.
@@
BEVformerV2[31] - Cam R50 52.9 42.3 0.618 0.273 0.413 0.333 0.181
BEVformerV2[31] - Cam InternImage-B 48.5 41.7 0.696 0.275 0.354 0.734 0.182
@@
CLIP-BEVFormer-tiny MLP Cam R50 38.5 27.4 0.869 0.283 0.607 0.542 0.220
CLIP-BEVFormer-tiny LM Cam R50 38.8 27.3 0.856 0.282 0.583 0.538 0.228
@@
CLIP-BEVFormer-base MLP Cam R101 56.2 46.7 0.605 0.253 0.331 0.336 0.187
CLIP-BEVFormer-base LM Cam R101 55.1 44.1 0.641 0.253 0.319 0.307 0.172
@@ 6 Experiment Setup
7 3D Object Detection Results on Test Set
The 3D object detection results on the test set are presented in Tab. 6. Notably, our proposed CLIP-BEVFormer consistently outperforms the baselines, showcasing significant improvements. Specifically, for the tiny-variant, we achieve a notable increase of 10.5% and 7.3% in NDS and mAP, respectively. In the case of the base-variant, our model demonstrates improvements of 2.2% in NDS and 0.5% in mAP.
@@
BEVFormer-tiny[17] - C R50 37.2 27.3 0.856 0.283 0.609 0.753 0.146
CLIP-BEVFormer-tiny LM C R50 41.1 29.3 0.811 0.271 0.554 0.579 0.136
BEVFormer-base[17] - C R101 53.5 44.5 0.631 0.257 0.405 0.435 0.143
CLIP-BEVFormer-base LM C R101 54.7 44.7 0.591 0.257 0.417 0.371 0.128
Table 6: 3D detection results on nuScenes test set.

### v2

@@ 1 Introduction
Autonomous Driving Systems (ADS) have witnessed rapid advancements, revolutionizing the landscape of transportation [13, 9, 14, 16, 6, 19, 20, 21]. The current state-of-the-art vision based ADS approaches heavily rely on Bird’s Eye View (BEV) representations extracted from multi-view images to perceive and understand the surrounding environment. The BEV detector, which acts as the backbone of ADS, transforms multi-view images into a top-down view representation known as the BEV feature map [24, 22, 12, 17, 36]. The effectiveness of the BEV detector significantly influences the success of perception tasks in autonomous driving, underscoring its pivotal role in enhancing the system’s situational awareness [17, 14, 16].
@@ 2.2 Vision-Language Models
Vision-language models (VLMs) have shown great promise for learning good representations for variety of downstream tasks [26, 27, 41, 39, 15, 1, 23]. The success of VLMs have been driven by training transformers on large scale image-text pairs data collected from web using contrastive learning  [26, 15]. Among VLMs, CLIP [26] which have been trained on 400 million pairs of data have shown remarkable zero-shot generalization on various image recognition tasks. The contrastive learning mechanism which maps image and text pairs to a joint embedding space has been pivotal for CLIP success. While VLMs generally show good performance on downstream tasks without any fine-tuning, prompting and fine-tuning techniques have also been employed for adaptation of VLMs to a new downstream task  [45, 10].
@@ 4.1 Implementation Details
Dataset. Our experiments are conducted on public nuScenes dataset[3], which is a commonly used dataset in autonomous driving area. It contains 1000 scenes, each with a $\sim$20s duration, and the keyframes are annotated at 2Hz. Each scene is captured with 6 cameras covering the entire 360 field-of-view. Overall, the nuScenes dataset contains 1.4M 3D annotated bounding boxes of 10 object categories.
@@
BEVerse [42] - Cam Swin 46.6 32.1 0.681 0.278 0.466 0.328 0.190
+Ours LM Cam Swin 48.3 34.2 0.665 0.270 0.456 0.318 0.170
@@
+Ours MLP Cam R50 38.5 27.4 0.869 0.283 0.607 0.542 0.220
+Ours LM Cam R50 38.8 27.3 0.856 0.282 0.583 0.538 0.228
@@
+Ours MLP Cam R101 56.2 46.7 0.605 0.253 0.331 0.336 0.187
+Ours LM Cam R101 55.1 44.1 0.641 0.253 0.319 0.307 0.172
BEVformerV2 [35] - Cam R50 42.6 35.1 0.753 0.286 0.466 0.807 0.186
+Ours LM Cam R50 44.1 37.0 0.729 0.281 0.438 0.791 0.204
@@ 6 Experiment Setup
7 3D Object Detection Results with Various Baselines
We have conducted experiments with various detection baselines, BEVformer [17], BEVformerV2 [35], and BEVerse [42]. We evaluate our model on both validation and test sets of nuScenes. The results presented in Tab. 6 show that our method consistently improves the perception capabilities of various baselines by significant margins on both sets, indicating its flexibility and model-agnostic nature.
Model Backbone Validation Set Test Set
@@
BEVformer-tiny R50 35.5 25.1 0.898 0.293 0.651 0.657 0.216 37.2 27.3 0.856 0.283 0.609 0.753 0.146
+Ours R50 38.8 27.3 0.856 0.282 0.583 0.538 0.228 41.1 29.3 0.811 0.271 0.554 0.579 0.136
BEVformer-base R101 51.7 41.6 0.673 0.274 0.372 0.394 0.198 53.5 44.5 0.631 0.257 0.405 0.435 0.143
+Ours R101 55.1 44.1 0.641 0.253 0.319 0.307 0.172 54.7 44.7 0.591 0.257 0.417 0.371 0.128
BEVformerV2 R50 42.6 35.1 0.753 0.286 0.466 0.807 0.186 42.5 35.4 0.707 0.278 0.506 0.895 0.134
+Ours R50 44.1 37.0 0.729 0.281 0.438 0.791 0.204 43.6 37.9 0.676 0.272 0.475 0.975 0.141
BEVerse Swin 46.6 32.1 0.681 0.278 0.466 0.328 0.190 50.1 36.2 0.610 0.257 0.451 0.355 0.131
+Ours Swin 48.3 34.2 0.665 0.270 0.456 0.318 0.170 52.2 37.4 0.556 0.247 0.413 0.301 0.129
Table 6: 3D object detection results on nuScenes validation and test sets.
@@
9 Training and Inference Efficiency
Our model is trained with 4 A100 80GB GPUs. Our method does not introduce any additional parameters and computations during the inference stage, which means that it allows for enhanced performance without sacrificing real-time processing capabilities. We provide detailed information on memory, training time, number of parameters and FPS in Tab. 7.
Model Train Mem (GB) Train Hrs # Params (M) FPS
BEVformer-tiny $\sim$7 $\sim$46 33 5.1
+Ours $\sim$7 $\sim$46 33 5.1
BEVformer-base $\sim$25 $\sim$90 69 2.1
+Ours $\sim$25 $\sim$90 69 2.1
BEVformerV2 $\sim$46 $\sim$38 56 2.3
+Ours $\sim$46 $\sim$38 56 2.3
BEVerse $\sim$48 $\sim$72 102.5 4.4
+Ours $\sim$48 $\sim$72 102.5 4.4
Table 7: Efficiency details. FPS is tested on 1 V100 GPU.